### v1

# Error thresholds of toric codes with transversal logical gates

Abstract
The threshold theorem promises a path to fault-tolerant quantum computation by suppressing logical errors, provided the physical error rate is below a critical threshold. While transversal gates offer an efficient method for implementing logical operations, they risk spreading errors and potentially lowering this threshold compared to a static quantum memory. Available threshold estimates for transversal circuits are empirically obtained and limited to specific, sub-optimal decoders. To establish rigorous bounds on the negative impact of error spreading by the transversal gates, we generalize the statistical mechanical (stat-mech) mapping from quantum memories to logical circuits. We establish a mapping for two toric code blocks that undergo a transversal CNOT (tCNOT) gate. Using this mapping, we quantify the impact of two independent error-spreading mechanisms: the spread of physical bit-flip errors and the spread of syndrome errors. In the former case, the stat-mech model is a 2D random Ashkin-Teller model. We use numerical simulation to show that the tCNOT gate reduces the optimal bit-flip error threshold to $p=0.080$, a $26\%$ decrease from the toric code memory threshold $p=0.109$. The case of syndrome error coexisting with bit-flip errors is mapped to a 3D random 4-body Ising model with a plane defect. There, we obtain a conservative estimate error threshold of $p=0.028$, implying an even more modest reduction due to the spread of the syndrome error compared to the memory threshold $p=0.033$. Our work establishes that an arbitrary transversal Clifford logical circuit can be mapped to a stat-mech model, and a rigorous threshold can be obtained correspondingly.
I Introduction
The threshold theorem [1, 2, 3, 4] is a cornerstone of quantum error correction (QEC), promising a path toward fault-tolerant quantum computation (FTQC) [5, 6, 7]. The theorem states that for a quantum device with a physical error rate below a non-zero threshold, a QEC scheme can suppress the logical error rate to an arbitrarily low level. For instance, the toric code used as a quantum memory has a well-established bit-flip error threshold of approximately 10.9% for perfect syndrome measurements, or 3.3% when including syndrome errors [8, 9, 10]. Encouragingly, recent hardware breakthroughs have pushed various quantum platforms into this sub-threshold regime, enabling demonstrations where error correction improves the fidelity of logical memory [11, 12]. These advances set the stage for the implementation of fault-tolerant logical computations [13, 14, 15, 16, 17]. Two dominant schemes for this task are transversal gates [18, 19, 20] and lattice surgery [21], each having its own pros and cons. Specifically, compared to lattice surgery, transversal gates require fewer physical qubits and a shallower circuit to implement. However, physical errors propagate across code blocks via physical entangling operations under the transversal gate. It is to be expected, therefore, that the presence of transversal gates will lower the error threshold. Hence, rigorously establishing how transversal gates affect the error threshold has vast implications for the pursuit of fault-tolerant computation.
Although Ref. [22] proved the existence of a non-zero error threshold of transversal logical circuit, the actual error threshold has only been observed empirically for specific, practical decoders [23, 24, 25, 26, 27, 28, 29, 30]. Therefore, a decoder-agnostic maximal error threshold is much needed for transversal logic circuits. For quantum memories, these tasks are accomplished through the statistical mechanical (stat-mech) mapping [8, 9], which maps the physical error rate to temperature and disorder strength, and the threshold to the critical point of order-disorder phase transition. The mapping leverages the correspondence between the large code distance limit and the thermodynamic limit of the statistical mechanics model to establish a rigorous threshold. However, to this date, stat-mech models in the presence of transversal gates are not known. In fact, rigorous derivation of error threshold using stat-mech mapping are only available for static quantum memories (see Table 1 for a list of existing stat-mech mappings for QEC codes).
In principle, the maximal error threshold of transversal logical circuits can also be crudely estimated from circuit simulation with a most likely error (MLE) decoder [24, 22], due to its near-optimal performance for toric code. However, since finding the MLE in general is an NP-hard problem for logical circuits, such an empirical approach is limited to small code distances. By deriving a mapping between transversal logical circuit under error and a classical stat-mech model, and simulating the model with finite size scaling to infer the critical point in the thermodynamic limit, we aim to establish a rigorous threshold in the infinite code distance limit (thermodynamic limit of stat-mech model).
In this work, we show that the logical circuit applying a tCNOT gate to two toric code blocks with persistent bit-flip errors and perfect syndromes maps to a random Ashkin-Teller (AT) model. We perform classical Monte-Carlo simulations of the random AT model, which shows that the gate lowers the optimal bit-flip error threshold of the target code block to $p=0.080$, corresponding to a $26\%$ decrease from the memory threshold. With syndrome errors, we show that the logical circuit maps to a 3D random 4-body Ising model with a plane defect inserted in the time slice when the gate is applied. We then argue, using reported numerical results on 3D random 4-body Ising model, that the spreading of syndrome errors causes a more modest decrease of at most 14% in the target block’s threshold, compared to the memory threshold.
This paper is organized as follows. In Sec. II, we review basics of stat-mech mapping for toric code quantum memory. In Sec. III, we introduce the tCNOT gate and spacetime detectors across the tCNOT gate. We then define the first noise models of our interest, namely the persistent bit-flip error and derive the stat-mech model of the tCNOT gate with perfect syndromes. We conduct Monte-Carlo simulation of the stat-mech model and obtain the corresponding error threshold in this case. In Sec. IV, we derive the stat-mech model for the second noise model, where the bit-flip error is combined with errors in syndrome extractions, and estimate the reduction of error threshold due to the tCNOT gate. Finally, in Sec. V, we discuss the stat-mech model of a general transversal logical circuit and its implication in fault-tolerant quantum computations.
Logical ops? Stabilizer code Error model syndrome error? Stat-mech model
No 2D toric code Bit/phase flip No 2D RBIM($p_{\text{th}}\approx 0.109$) [8]
Yes 3D R4bIM($p^{*}\approx 0.033$) [9, 10]
Depolarizing No 2D Random AT model [31]
Correlated error No RBIM with further neighbor term [32]
Single-qubit coherent error No Random complex coupling AT model [33, 34]
Both coherent and incoherent bit flip No
2D color code Bit/phase flip No 2D R3bIM [35, 36]
Yes 3D Random Ising gauge theory[37]
3D toric code Bit/phase flip No 3D RBIM/R4bIM [10, 38] [^1] The 3D toric code and color code are not self-dual, hence the stat-mech models of bit and phase flip errors are different.
3D color code Bit/phase flip No 3D R4bIM/R6bIM [39]
Yes 2D toric code + tCNOT (this work) Persistent bit/phase flip No 2D Random AT model
bit/phase flip Yes 3D R4bIM + plane defect
Table 1: Existing stat-mech mapping for 2D and 3D QEC codes. RBIM: random link Ising model. AT: Ashkin-Teller. RnnbIM ($n=3,4,6$): random $n$-body Ising model. 3D R4bIM is also known in the literature as random plaquette Ising gauge theory.
II Stat-mech mapping for error threshold of quantum memories
Determining the error threshold using stat-mech mapping consists of two steps. The first step is to map the probability distribution of syndromes to the partition function of a stat-mech model. The second step is to identify the thermodynamic quantity in the stat-mech model that corresponds to the logical error rate. Since our derivation closely follows the derivations of quantum memories, we first review the stat-mech mapping for quantum memories for completeness, following Refs. [8, 9, 32].
Figure 1: Stat-mech mapping of toric code under errors. (a) Sketch of trivial $C\in S$ and non-trivial cycle $C\in L$ on top of an error $E$ in one toric code block, which triggers a pair of syndromes denoted by the two red dots. (b) Stat-mech model of toric code under bit-flip errors. Here the physical qubits of the toric code are marked by circles, and the weight-4 Z stabilizers are defined on the shaded plaquettes. A trivial cycle $C\in S$ can be parameterized by a domain wall of Ising spins $\{\sigma\}$ (marked by the arrows) on a 2D square lattice that is rotated by $45^{\circ}$, marked by dashed lines. Two ends of a link $l$ on this lattice are denoted by $l_{1,2}$. (c) The 3D stat-mech model in the case of syndrome errors for a single code block. The Ising spins live at the links of thd 3D cubic lattice, marked by the arrows. Some of the 4-body interaction terms in the Hamiltonian in Eq. (II.2) are marked by grey stars. The dashed lines at $t\in\mathbb{Z}+\frac{1}{2}$ form a 2D square lattice where Ising spins live on the sites, which is the same 2D square lattice in (b).
II.1 Stat-mech mapping with perfect syndromes
We first consider the case with perfect syndromes, where every stabilizer measurement returns the actual physical value of the stabilizer. In this case, the error that occurs in the code state consists solely of physical Pauli errors, which will be denoted as a Pauli operator $E$. The error model is defined as the probability distribution $\mathrm{prob}(E)$. Physically, the Pauli error is a quantum channel that maps the density matrix of the system $\rho$ to $\sum_{E}\mathrm{prob}(E)E\rho E^{\dagger}$. In this case, the logical error rate, $P_{\text{logical}}$, can be computed as follows [8]:
$P_{\text{logical}}$ $=\sum_{E}\mathrm{prob}(E)\left[\sum_{C\in L}\mathrm{prob}(E\cdot C|E)\right]$
$=\sum_{E}\mathrm{prob}(E)\frac{\sum_{C\in L}\mathrm{prob}(E\cdot C)}{\sum_{C\in S\cup L}\mathrm{prob}(E\cdot C)},$ (1)
where $C$ denotes a “cycle” of Pauli operators that do not trigger any syndromes when the stabilizers are measured [^2] We call $C$ a cycle since in 2D topological codes such an operator is usually supported on closed cycles.. That is, it either belongs to the stabilizer group of the code state, denoted by $S$, or the set of non-trivial logical operator, denoted by $L$. In this way, Pauli errors $E$ and $E\cdot C$ will trigger exactly the same set of syndromes. In the first equality, the expression in square brackets is the probability of logical error given that a physical error $E$ occurs, which is computed via the second equality. Without the knowledge of the ground truth about $E$, the optimal maximum likelihood decoder guesses an error $E\cdot C$ according to the conditional probability $\mathrm{prob}(E\cdot C|E)$. A logical error will occur whenever our guess differs from the actual physical error by any logical operator, see Fig. 1(a). Averaging over all possible errors yields the logical error rate in Eq. (II.1).
The central step of stat-mech mapping is to rewrite the summation of the cycles $C$ in Eq. (II.1) into a partition function of a stat-mech model with quenched disorder. For toric code under single qubit bit flip error, the cycle $C$ can be parameterized using a set of Ising spins $\{\sigma\}$ living on the lattice sites of a 2D square lattice, see Fig. 1(b). Topologically trivial cycle $C\in S$ is mapped to a topologically trivial domain wall of Ising spins. The summation of $C\in S$ now becomes the following partition function:
$\sum_{C\in S}\mathrm{prob}(E\cdot C)\propto Z(E)=\sum_{\{\sigma\}}e^{-H(\{\sigma\}|E)},$
$H(\{\sigma\}|E)=-J\sum_{l}s_{l}\sigma_{l_{1}}\sigma_{l_{2}},$ (2)
where $H(\{\sigma\}|E)$ is the Hamiltonian of the 2D random bond Ising model (RBIM). Here $J=\frac{1}{2}\ln\frac{1-p}{p}$ is related to the bit-flip error rate $p$, $l_{1,2}$ are two lattice sites at the ends of the link $l$, and the sign of the random bond interaction, $s_{l}=1/-1$, corresponds to the absence/existence of physical errors of the qubit that is located at the link $l$.
Meanwhile, the sum over the non-trivial cycles $C\in L$ becomes the sum of partition functions with all possible non-trivial global domain walls (DW) in the Hamiltonian, i.e.
$\sum_{C\in L_{a}}\mathrm{prob}(E\cdot C)\propto Z(E)\equiv\sum_{\{\sigma\}}e^{-H_{DW_{a}}(\{\sigma\}|E)},$ (3)
where $a=1,2,3$ denotes the three topologically non-trivial global domain wall configurations around the 2D torus, each giving rise to a different logical error. Therefore, the logical error rate in terms of the stat-mech model is
$P_{\text{logical}}=\sum_{E}\mathrm{prob}(E)\sum_{a}e^{-F_{DW_{a}}(E)}=\sum_{a}\left\langle{e^{-F_{DW_{a}}}}\right\rangle,$ (4)
where $F_{DW_{a}}(E)\equiv-\ln\frac{\sum_{\{\sigma\}}e^{-H_{DW_{a}}(\{\sigma\}|E)}}{\sum_{\{\sigma\}}e^{-H(\{\sigma\}|E)}}$ is the domain wall free energy of a global domain wall in the topological class $a$, and $\langle\cdots\rangle$ denotes the disorder average over the error configuration $E$.
The error threshold now has a clear physical picture: the critical point of the order-disorder phase transition of the stat-mech model. For low error rate, the stat-mech model has a low effective temperature (i.e. the couplings in $H(\{\sigma\}|E)$ are strong), and the stat-mech model is in the ordered phase. Consequently, the domain wall free energy $F_{DW}$ scales with the system size. Therefore, the logical error rate is exponentially suppressed with increasing code distance. In contrast, the high error rate is mapped to the disordered phase of the stat-mech model at high temperature, where the domain wall free energy $F_{DW}$ is finite. Therefore, $P_{\text{logical}}$ remains finite even in the large code distance limit. Based on numerical simulations in Ref. [40], the error threshold is $p_{\text{th}}=0.109$.
II.2 Stat-mech mapping with syndrome errors
We then review the stat-mech mapping for the case of syndrome errors. A common way to model this is to assume that every stabilizer measurement has the probability $q$ being the opposite of its true physical value. In this case, one generally needs multiple rounds of syndrome extractions to suppress logical errors. Therefore, the corresponding stat-mech model becomes three dimensional [9]. Again, the core in the stat-mech mapping procedure is to represent the summations over $C$ in Eq. (II.1) in partition functions. In the case of toric code memory under bit-flip noise with syndrome errors, every cycle $C$ can be parameterized by a set of Ising spins $\{\sigma\}$, which lives on the links of the 3D cubic lattice, see Fig. 1(c). The corresponding stat-mech model is the 3D random 4-body Ising model(R4bIM), also known as the random plaquette Ising gauge theory [8, 9, 10]:
$\sum_{C}\mathrm{prob}(E\cdot C)\propto Z_{3D}(E)=\sum_{\{\sigma\}}e^{-H_{3D}(\{\sigma\}|E)},$
$H_{3D}(\{\sigma\}|E)=-K\sum_{p,t}r_{p}(t)\prod_{l\in p}\sigma_{l}(t)$
$-J\sum_{l,t}s_{l}(t+\frac{1}{2})\sigma_{l}(t)\sigma_{l}(t+1)\sigma_{l_{1}}(t+\frac{1}{2})\sigma_{l_{2}}(t+\frac{1}{2}).$ (5)
Here, the coupling constants are related to the error rates via $J=\frac{1}{2}\ln\frac{1-p}{p}$ and $K=\frac{1}{2}\ln\frac{1-q}{q}$. The error $E$ is parameterized by two sets of binary numbers: $s_{l}(t+\frac{1}{2})=1/-1$ for the absence/occurrence of a physical error at the data qubit on the link $l$, and $r_{p}=1/-1$ for the absence/occurrence of a syndrome error of the $Z$ stabilizer at the plaquette $p$.
In terms of the stat-mech model, the bit-flip error threshold of toric code with $O(d)$ rounds of syndrome extraction is mapped to the confinement transition of the R4bIM at $d\to\infty$ with increasing $p$ and $q$. At low error rates, the stat-mech model is in the deconfined (Higgs) phase, signaled by the perimeter-law decay of the Wilson loop order parameter $\langle W(\gamma)\rangle\equiv\langle\prod_{l\in\gamma}\sigma_{l}\rangle\sim e^{-A|\gamma|}$, where $\gamma$ is a closed loop in the cubic lattice, $|\gamma|$ is its length, and $A>0$ is some constant. Here, $\langle\dots\rangle$ represents both averaging over the partition function $Z_{3D}(E)$ and then averaging over the disorder $E$. At high error rates, the model enters the confined phase where the expectation value of the Wilson loop $W(\gamma)$ decays with the area $\mathfrak{S}_{\gamma}$ that the loop $\gamma$ encircles: $\langle W(\gamma)\rangle\sim e^{-B\mathfrak{S}_{\gamma}}$, where $B>0$ is some constant. The critical point of the confinement transition is around $p^{*}\approx 0.033$ in the case of $p=q$ [10].
III Toric codes under tCNOT gate with perfect syndromes
We now generalize the stat-mech mapping to logical circuits involving the tCNOT gate between two toric code blocks. Physically, the tCNOT gate is implemented via physical CNOT gates between every pair of physical qubits (see Fig. 2(a)). Since bit-flip operation on the control qubit does not commute with CNOT gate, the error will spread. In particular, we consider a specific bit-flip error model in this circuit that manifests the error spreading of bit-flip error by the tCNOT gate, and study its effect on the error threshold. The logical error rate in this case will still have the form of Eq. (II.1). However, we now need to determine the form of the cycle $C$ in logical circuits, in order to sum over all the cycles. The cycle $C$ is still defined as a set of errors, such that for any error $E$, $E$ and $E\cdot C$ trigger the same set of syndromes. The major difference here is that the notion of syndrome needs to be generalized to incorporate the propagation of errors by the transversal logical gates. Therefore, we first introduce the bit-flip error model of our interest, and then review error syndromes across a tCNOT gate.
III.1 Persistent bit-flip error model
For quantum memories, bit-flip errors between two consecutive rounds of syndrome extractions are modeled via the following quantum channel:
$\mathcal{N}_{p}\equiv\circ_{i}\mathcal{N}_{p,i},\ \mathcal{N}_{p,i}(\rho)\equiv(1-p)\rho+pX_{i}\rho X_{i}.$ (6)
Here, $p$ is the cumulative probability that a physical bit-flip error occurs between two consecutive rounds of syndrome extraction, and $i$ labels different physical qubits in the system. However, since bit-flip operation on the control qubit does not commute with the physical CNOT gate, the error model in a logical circuit needs to further specify the temporal location of the bit-flip errors.
Therefore, we consider an error model that captures the spreading of error, where bit-flip errors occur both before and after the tCNOT gate. To directly compare with the case of quantum memory, we evenly split the error rate of channel $\mathcal{N}_{p}$ into two bit-flip channels $\mathcal{N}_{\tilde{p}}$ before and after the tCNOT gate:
$\mathcal{N}_{\tilde{p}}\circ\mathcal{N}_{\tilde{p}}=\mathcal{N}_{p},$ (7)
as illustrated in Fig. 2(b). In order for the combined error channels to have a net bit-flip probability $p$, we must have
[…]
We will subsequently refer to such an error model as the persistent bit-flip error [^3] When referring to the bit-flip error without the word “persistent”, we assume the bit-flip channel in Eq. (6) occurs only after the tCNOT gate, which is the case in the noise model we study in Sec. IV. .
To make a direct comparison with the memory thresholds, we assume that every physical CNOT gate in the tCNOT gate is free from noise. In reality, the errors of the physical CNOT gates can be incorporated into the error channels $\mathcal{N}_{\tilde{p}}$ before and after them.
III.2 Detectors of toric codes across a tCNOT gate
The main challenge in applying stat-mech mapping to logical circuits lies in the necessity of tracking dynamical error propagation through gates. For this, we import the quantum circuit simulation notion of detector [41], and integrate it into the stat-mech mapping procedure. In our case, a detector associated with a stabilizer measurement outcome is a product of it with earlier stabilizer measurement outcomes that back-tracks the flow of that measured Pauli operator in the circuit. Without errors, the value of each detector is deterministically 1. When an error $E$ occurs in the circuit, a detector can be triggered, which flips its value to -1. The collection of detector values, $\{d\}$, serves as the syndrome. Therefore, once the detectors and its relation with every error $E$ in the error model is specified, so will the cycles $C$.
In the context of a logical circuit with a tCNOT gate, the spacetime detectors are fixed in the following way [24, 26, 27, 28, 29]. Since we only consider bit-flip errors, we back-track the spread of the measured $Z$ stabilizers. In this way, the spacetime detectors between the two adjacent syndrome extraction rounds are the ones shown in Fig. 2(a): the spacetime detector $d^{c}$ of the control code block is the product of the measurement outcomes $M^{c}_{1/2}$ of the $Z$ stabilizers before/after the tCNOT gate:
$d^{c}=M^{c}_{1}M^{c}_{2}.$ (9)
Meanwhile, the spacetime detector $d^{t}$ of the target code block to be the product is now
$d^{t}=M^{c}_{1}M^{t}_{1}M^{t}_{2},$ (10)
where $M^{t}_{1,2}$ are outcomes of the $Z$ stabilizer measurements before and after the tCNOT gate. The superscript denotes the control code ($c$) and the target code ($t$) respectively (alternatively, one can also choose $d^{t}=M^{c}_{2}M^{t}_{1}M^{t}_{2}$.). A visualization of the detectors is show in Fig. 2(a). The spacetime detectors directly incorporate both error propagation mechanisms caused by the tCNOT gate. Namely, a bit-flip error that occurs in the error channel $\mathcal{N}_{\tilde{p}}$ before the tCNOT gate in the control block now triggers both $d^{c}$ and $d^{t}$.
Figure 2: (a) The implementation of the tCNOT gate between two rounds of syndrome extractions. The black dots at the centers of shaded plaquettes represent the measurement of the $Z$ stabilizers defined on these plaquettes. The choices of spacetime detectors $d^{c}$ and $d^{t}$ for the control and target code blocks between the two rounds of SE are marked in blue and green, repectively. Due to the tCNOT gate, the spacetime detector of the target block, $d^{t}$, extends to the control block, whose value is obtained from multiplying three $Z$ stabilizer measurement outcomes $d^{t}=M^{c}_{1}M^{t}_{1}M^{t}_{2}$. To avoid visual overlap, we sketch the spacetime detectors $d^{c}$ and $d^{t}$ in two adjacent plaquettes that host $Z$ stabilizers. (b) The logical circuit with tCNOT gate, persistent bit-flip noise channels $\mathcal{N}_{\tilde{p}}$ and two rounds of perfect syndrome extractions (SE), during which the weight-4 $Z$ stabilizers are measured.
III.3 Stat-mech mapping for perfect syndromes and persistent bit-flip errors
With the detectors and error model fixed, we now construct the stat-mech model of the two toric code blocks across the tCNOT gate. We denote the bit-flip errors that occur during the four error channels $\mathcal{N}_{\tilde{p}}$ in Fig. 2(a) as $E^{c,t}_{1,2}$, where $c/t$ indicates the control/target code blocks and $1/2$ denotes the error before/after the tCNOT gate. Given the form of the spacetime detectors $d^{c/t}$, the errors that trigger spacetime detectors in the control and target blocks are
$E^{c}=E^{c}_{1}\cdot E^{c}_{2}\text{ and }E^{t}=E^{t}_{1}\cdot E^{t}_{2}\cdot E^{c}_{1},$ (11)
respectively. Since every physical bit-flip error occurs independently, the probability $\mathrm{prob}(E^{c},E^{t})$ is a product of independent probability distributions between pairs of physical qubits in the two code blocks that undergo physical CNOT gates:
$\mathrm{prob}(E^{c},E^{t})=\prod_{l}p(s^{c}_{l},s^{t}_{l}),$ (12)
where $l$ is the index of the spatial location of the physical qubits, which now on the links of a rotated square lattice (marked in dotted lines in Fig. 3), $s_{1,2}^{c,t}=1/-1$ to denote the absence/occurrence of a bit-flip in each respective error channels, and $p(s^{c}_{l},s^{t}_{l})$ denotes the joint probability distribution of bit-flip errors occurring to the two physical qubits at $l$.
We focus on a pair of physical qubits in the control and target block. Since the error channels $\mathcal{N}_{\tilde{p}}$ are identical, the errors in the two channels follow the same probability distribution:
$p(s_{1,2}^{c,t})=\frac{1+(1-2\tilde{p})s^{c,t}_{1,2}}{2}.$ (13)
Denoting the cases where the combined errors will not trigger/trigger the control and target spacetime detectors as $s^{c,t}=1/-1$, we have
$p(s^{c},s^{t})=$ $\sum_{s_{1,2}^{c,t}=\pm 1}p(s_{1}^{c})p(s_{2}^{c})p(s_{1}^{t})p(s_{2}^{t})$
$\quad\times\delta(s^{c}-s_{1}^{c}s_{2}^{c})\delta(s^{t}-s^{t}_{1}s^{t}_{2}s^{c}_{1})$
$=$ $\frac{1+(1-2\tilde{p})^{2}s^{c}+(1-2\tilde{p})^{3}s^{t}(1+s^{c})}{4}.$ (14)
In the first equality, the delta functions enforce the two conditions in Eq. (11), while the second equality carries over the summation using Eq. (13). The joint distribution cannot be factorized into a product of individual probability distributions, which means $s^{c}$ and $s^{t}$ are correlated random variables. This is exactly due to the shared error $E^{c}_{1}$ between $E^{c}$ and $E^{t}$ in Eq. (11).
With the representation of errors and the spacetime detectors $d^{c,t}$, we can now find a parameterization of the cycle $C$. With perfect syndromes, the spacetime detector on the plaquette $p$ in the control/target code block is triggered depending on the sign of $d^{c/t}_{p}=\prod_{l\in p}s^{c/t}_{l}$. Therefore, the following transformation of the physical errors will preserve the value of $d^{c/t}_{p}$:
$s^{c}_{l}\to s^{c}_{l}\sigma_{l_{1}}\sigma_{l_{2}},\ s^{t}_{l}\to s^{t}_{l}\tau_{l_{1}}\tau_{l_{2}},$ (15)
where $\{\sigma=\pm 1\},\{\tau=\pm 1\}$ are two sets of Ising spins that live on the lattice sites, and $l_{1},l_{2}$ are the two lattice sites at the two ends of the link $l$, similar to the labeling in Fig. 1(b). In this way, the trivial cycle $C$ can be separated into two parts: $C^{c}$ for the control block parameterized by $\{\sigma\}$, and $C^{t}$ for target code block parameterized by $\{\tau\}$, see Fig. 3. The stat-mech model can then be obtained via rewriting the summation of cycles using the Ising spins:
$\sum_{C^{c}\in S^{c},C^{t}\in S^{t}}\mathrm{prob}(E^{c}\cdot C^{c},E^{t}\cdot C^{t})$
$=\sum_{\{\sigma\},\{\tau\}}\prod_{l}p(s^{c}_{l}\sigma_{l_{1}}\sigma_{l_{2}},s^{t}_{l}\tau_{l_{1}}\tau_{l_{2}})$
$\propto Z(E^{c},E^{t})=\sum_{\{\sigma\},\{\tau\}}e^{-H({\{\sigma\},\{\tau\}}|E^{c},E^{t})},$ (16)
where $S^{c,t}$ denote the stabilizers in the control and target code blocks, respectively. The disordered Hamiltonian is a 2D Ashkin-Teller (AT) model on the square lattice with random link couplings [42, 43, 44, 45, 46, 47]:
$H({\{\sigma\},\{\tau\}}|E^{c},E^{t})=-K_{2}\sum_{l}s^{c}_{l}\sigma_{l_{1}}\sigma_{l_{2}}$ (17)
$\quad\quad-K_{4}\sum_{l}s^{t}_{l}\tau_{l_{1}}\tau_{l_{2}}-K_{4}\sum_{l}s^{t}_{l}s^{c}_{l}\tau_{l_{1}}\tau_{l_{2}}\sigma_{l_{1}}\sigma_{l_{2}},$
where the coupling constants $K_{2}$ and $K_{4}$ are functions of $\tilde{p}$ (see Appendix A for their explicit forms), and $s^{c,t}_{l}$ are quenched random variables following the probability distribution in Eq. (III.3). We note that random AT model also arises from stat-mech mapping of toric code quantum memory under depolarizing noise [31] and both coherent and incoherent bit-flip noises [48, 34]. Our model in Eq. (17) differs from the existing models in the coupling constants and disorder distributions. Compared to RBIM for the memory case in Eq. (II.1), the two coupling constants in the random AT model satisfy
$K_{4}<K_{2}<J=\frac{1}{2}\ln\frac{1-2\tilde{p}(1-\tilde{p})}{2\tilde{p}(1-\tilde{p})}$ (18)
for $\tilde{p}$ around the memory threshold (see Fig. 4). The relations between these coupling strengths have direct implications on the error thresholds of the control and target code blocks, as shown in the subsequent discussion.
Figure 3: Stat-mech mapping of two toric code blocks undergoing a tCNOT gate, where the probability distribution of $E^{c}$ and $E^{t}$ are correlated due to the tCNOT gate. Here the physical qubits are denoted by gray circles living on the lattice makred by solid lines. Bit-flip errors $E^{c/t}$ and $E^{c/t}\cdot C^{c/t}$ of the physical qubits in the control/target code blocks are marked by gray circles with colored fillings. Syndromes triggered by $E^{c/t}$ are marked by dots at the center of shaded plaquettes which host $Z$ stabilizers. Trivial cycles $C^{c/t}$ in control/target blocks are parameterized by two sets of Ising spins $\{\sigma\}$ and $\{\tau\}$, which are located at lattice sites of the dotted-line lattice, and are denoted by arrows in blue/green.
Figure 4: Coupling constants $K_{2}$ and $K_{4}$ in the random AT model Hamiltonian in Eq. (17), whose relations with $\tilde{p}$ are given in the Appendix Eq.(A). Coupling of RBIM, $J=\frac{1}{2}\ln\frac{1-2\tilde{p}(1-\tilde{p})}{2\tilde{p}(1-\tilde{p})}$, is shown as a reference.
The stat-mech model allows us to form a theoretical anticipation of the error thresholds of the two code blocks, even before carrying out the numerical simulation. In light of Eq. (4), the logical error rate for the control/target blocks now maps to the sum over $\overline{e^{-F_{DW}^{\sigma/\tau}}}$, where $F^{\sigma/\tau}_{DW}$ is the domain wall free energy of a global domain wall of the spins $\sigma/\tau$, and the disorder average is taken over both $E^{c}$ and $E^{t}$ (i.e. $s^{c}$ and $s^{t}$). Therefore, the error thresholds of the control and target code blocks, $p^{c}_{\text{c}}$ and $p^{t}_{\text{c}}$, correspond to the order-disorder phase transition of the $\sigma$ and $\tau$ spins, respectively. From Eq. (18), we expect that the target block will first reach the error threshold with increasing $\tilde{p}$. Since the critical point is within the ordered phase of $\{\sigma\}$ spins, we expect the coupling term between $\{\sigma\}$ and $\{\tau\}$ spins to enhance the error threshold of the target block $p^{t}_{\text{c}}$. On the other hand, since $K_{2}<J$, the error threshold of the control layer $p^{c}_{\text{c}}$ should also be reduced compared to the case of quantum memory, $p_{\text{th}}\approx 0.109$.
III.4 Numerical investigations of the error thresholds
III.4.1 Monte Carlo simulation of the disordered AT model
With the stat-mech model being constructed, we now conduct a numerical study to quantify the error thresholds of the two code blocks. Specifically, we perform a Monte Carlo (MC) simulation of the disordered AT model in Eq. (17), with disorders sampled from the distribution shown in Eq. (III.3). Unlike the conventional AT model, where $K_{2}$ and $K_{4}$ are independent parameters, our model depends on a single parameter $p$: $K_{2}$ and $K_{4}$ are both functions of $p$ (see Fig. 4). At each $p$, we first sample disorder configurations on all bonds $\{s_{l}^{t},s_{l}^{c}\}$ following Eq. (III.3). We then to equilibrate and sample the spin fields $\tau,\sigma$ with Hamiltonian described in Eq. (17) given quenched disorders $\{s_{l}^{t},s_{l}^{c}\}$. To ensure the MC simulation equilibrates, we introduce a hypothetical temperature $\beta$. The target distribution that we want to sample from is $e^{-\beta H}$ with $\beta=1$. We tune $\beta$ to effectively anneal the system from high temperature to target temperature $\beta=1$. We found this extra annealing equilibration stage to be crucial in reaching the equilibrium.
To observe the order-to-disorder transition, we analyze the magnetization of the spin system at different $p$. As shown in Fig. 5, the magnetization $\tau$ and $\sigma$ spin fields $|M|$ both saturate to $1$ in the low noise limit ($p=0$). Since we start the MC from a random initialization, $|M|=1$ in the low $p$ limit affirms the convergence of the simulation. As the noise level increases, $\tau$ magnetization, corresponding to target code block, first starts decreasing and eventually vanishes. The magnetization of $\sigma$, corresponding to the control code block, tolerates a larger noise level before it starts to deviate from $|M|=1$. This observation is consistent with the fact that bit-flip errors propagate from the control code block to the target code block through the tCNOT operation, resulting in a lower threshold for the target code than for the control code.
The major benefit of stat-mech mapping is the ability to leverage the established practice of finite-size scaling (FSS) to determine error thresholds in the infinite code distance limits. To obtain these thresholds, we need to collapse the magnetization of different system sizes $L=8,12,16$ into a unique scaling form [49]:
$ML^{\beta/\nu}=\mathcal{F}(\epsilon L^{1/\nu}),$ (19)
where $\epsilon=\frac{\tilde{p}-\tilde{p}_{\text{c}}}{\tilde{p}_{\text{c}}}$ is the rescaled distance from the critical point and $\mathcal{F}$ is some universal scaling function. We optimize three parameters $\beta,\nu,\tilde{p}^{\text{c}}$ to infer the critical error rate $p$ in the thermodynamic limit (infinite code distance).
As shown in Fig 6, we observe that the rescaled magnetization data achieve a remarkable collapse to universal functional forms around $\tilde{p}_{\text{c}}^{c}=0.052$ (or $p^{c}_{\text{c}}=0.099$, see Fig. 6(a)) and $\tilde{p}^{t}_{\text{c}}=0.042$ (or $p^{t}_{\text{c}}=0.080$, see Fig. 6(b)). These are compelling evidence of universal behavior in the vicinity of these critical points. Therefore, these critical points can be interpreted as upper bounds on decoding thresholds for control ($\sigma$) and target ($\tau$) code blocks in the large code distance limit. Compared to threshold of quantum memory, $p_{\text{th}}=0.109$,the thresholds of the control and the target blocks are lowered by $\frac{0.109-0.099}{0.109}=9.2\%$ and $\frac{0.109-0.080}{0.109}=26.6\%$, respectively. These results confirm our theoretical expectation at the end of Sec. III.3 regarding the hierarchy of the reduced thresholds. At the same time, our results establish that tCNOT gates do not lower the thresholds to a devastatingly degree.
Our finite size scaling results also highlight the benefit of correlated decoding. If one decodes the target block separately, the detectors of the target block effectively experience three consecutive bit-flip channels of $\mathcal{N}_{\tilde{p}}$. The error threshold of such a channel is the solution to
$\tilde{p}^{3}+3(1-\tilde{p})^{2}\tilde{p}=p_{\text{th}}=0.109,$ (20)
which is $\tilde{p}=0.039$ or $p=0.076$. We note that such a threshold is crudely estimated in Ref. [50] to be $\frac{2}{3}p_{\text{th}}=0.073$. Hence, the improvement of the error threshold due to joint decoding is $\frac{p^{t}_{c}-0.076}{0.076}=5.3\%$. In the stat-mech model, this improvement stems from the coupling term in the random AT model: since the $\{\tau\}$ spins’ critical point lies within the ordered phase of $\{\sigma\}$ spins, the coupling in Eq. (17) induces an extra Ising coupling between the $\{\tau\}$ spins.
Figure 5: Magnetization of (a) $\tau$ and (b) $\sigma$ ($|M_{\tau}|$ and $|M_{\sigma}|$) versus persistent noise strength $\tilde{p}$ for system sizes $L=8,12,16$.
Figure 6: Finite size collapse of magnetization for (a) $\tau$ and (b) $\sigma$ according to scaling formula Eq. 19. The collapse is obtained from the following choice of parameters: $\beta=0.252,\nu=1.8,\tilde{p}^{t}_{\text{c}}=0.042,$ and $\tilde{p}^{c}_{\text{c}}=0.052$.
Figure 7: Logical error rate of a surface under same noise model using then most likely error decoder. The crossing point of curves for different system sizes indicates a decoding threshold. The gray lines represent the phase transition point predicted by Monte Carlo simulation of the random AT model.
III.4.2 Direct simulation of error correction protocol
We cross-check the thresholds predicted by the MC simulation of the AT model with those observed in a practical error correction protocol. Since the error thresholds for a planar surface code and the toric code are expected to be the same [8], we opt to simulate a rotated surface code for convenience. We use stim package [41] to simulate and sample syndromes from the same logical circuit shown in Fig. 2(a) under the noise model described in Fig. 2(b). We then decode the sampled syndromes using the most likely error (MLE) decoder [24] and record the resulting logical error rates (LERs). Although the MLE decoder is not scalable and thus not favorable in practical QEC, it serves as a valuable tool for validating our theoretical predictions and identifying performance bounds. As shown in Fig. 7, the LERs of the control and target code with different code distances each cross at one particular error rate, which is the observed decoding threshold. The crossings of LER for both control and target codes match the critical points obtained from FSS analysis of magnetization data from MC simulation of stat-mech mapped disordered AT model. Below the decoding threshold, increasing code distance suppresses logical errors.
IV Toric codes under tCNOT gate with syndrome errors
Figure 8: A toric code logical circuit with tCNOT gate, repeated noisy SE (marked by shaded boxes) and bit-flip noise channel $\mathcal{N}_{p}$.
To model syndrome errors, we assume that each stabilizer measurement has a probability $q$ of being incorrect. In addition, we assume that physical bit-flip errors occur with probability $p$ only between the tCNOT gate and the next syndrome extraction (but not between the syndrome extraction and the tCNOT gate), see Fig. 8. This error model singles out the propagation of syndrome error by the tCNOT gates.
Figure 9: (a) The modified 5-body interaction term in the Hamiltonian in Eq. (22) for the $\{\tau\}$ spins due to the tCNOT gate at t=Tt=T.(b) Closed surfaces $\mathfrak{S}$ in the high temperature expansion of the terms involving the $\{\tau\}$ spins in the partition function $Z_{3D}(E^{c},E^{t})$. On the intersection between $\mathfrak{S}$ and the $t=T$ plane, the coupled terms between $\{\sigma\}$ and $\{\tau\}$ spins in Eq. (22) gives rise to a Wilson loop of $\{\sigma\}$ spins, marked by the blue loop.
We now construct the stat-mech model for repeated syndrome error extractions around the tCNOT gate. Using the expression of the detectors in Eqs. (9) and (10), the cycle $C$ can be parameterized using two sets of Ising spins, denoted by $\{\sigma\}$ and $\{\tau\}$, which live on the links of a 3D cubic lattice. Summing over all possible cycles, we arrive at the partition function of the following stat-mech model (see Appendix B for derivation):
$Z_{3D}(E^{c},E^{t})=\sum_{C}\mathrm{prob}((E^{c},E^{t})\cdot C)$
$\quad\quad\quad\quad\quad\ =\sum_{\{\sigma\},\{\tau\}}e^{-H_{3D}(\{\sigma\},\{\tau\}|E^{c},E^{t})},$
$H_{3D}(\{\sigma\},\{\tau\}|E^{c},E^{t})=H_{3D}(\{\sigma\}|E^{c})$
$\quad\quad\quad+H_{3D}^{t\leq T}(\{\tau\}|E^{t})+H_{3D}^{t>T}(\{\sigma\tau\}|E^{t}).$ (21)
Here, each individual Hamiltonian on the RHS is the one of the R4bIM given in Eq. (II.2), while $t\geq T$ in $H_{3D}^{t\leq T}$ means that this Hamiltonian only contains terms that are defined on the plaquettes at or before $t=T$, and similarly for $H_{3D}^{t>T}$. In fact, the Hamiltonian $H_{3D}(\{\sigma\},\{\tau\}|E^{c},E^{t})$ represents two decoupled R4bIMs with a spin permutation defect plane at $t=T$ that exchanges $\tau\leftrightarrow\tau\sigma$. Redefining $\tau$ to be $\tau\sigma$ at $t>T$ in the partition function further decouples the two flavors of spins there. As a result, they are only coupled via the following set of 5-body interaction terms between $t=T$ and $t=T+1$ (see Fig. 9(a)):
$-J\sum_{l}s^{t}_{l}(T+\frac{1}{2})\sigma_{l}(T)\tau_{l}(T)\tau_{l}(T+1)\tau_{l_{1}}(T+\frac{1}{2})\tau_{l_{2}}(T+\frac{1}{2}).$ (22)
Due to the coupling terms in Eq. (22), we expect the error threshold of the target to be modified near $t=T$. However, it is difficult to obtain a directly analytic statement on the error threshold, and a numerical simulation of the entire stat-mech model in Eq.(IV) is required. Nevertheless, techniques in statistical mechanics like high- and low temperature expansion often shed light towards critical points through its ability of connecting seemingly different stat-mech models [51]. This includes the renowned Kramers-Wannier duality in the 2D Ising model, where the critical point can be computed exactly by comparing the expressions of high- and low temperature expansion.
Following these inspirations, we perform a high temperature expansion of the terms involving the $\{\tau\}$ spins in the Hamiltonian, with the assumption that the $\{\sigma\}$ spins are in the deconfined (decodable) phase when the $\{\tau\}$ spins reach the critical point. In this way, the partition function in Eq. (IV) can be written as
$Z_{3D}(E^{c},E^{t})\propto$
$Z_{c}(E^{c})\sum_{\mathfrak{S}}\left[4\tanh(J,K)\right]^{|\mathfrak{S}|}\prod_{p\in\mathfrak{S}}(s^{t}_{p},r^{t}_{p})\overline{\prod_{l\in\mathfrak{S}\cap(t=T)}\sigma_{l}}.$ (23)
Here, the summation above is over closed surface $\mathfrak{S}$ on the 3D cubic lattice. $\tanh^{|\mathfrak{S}|}(J,K)$ is a shorthand notation that represents the product of $|\mathfrak{S}|$ factors of $\tanh J$ and $\tanh K$. The factor $4$ comes from the fact that each classical spin is shared by two plaquettes, therefore summing over them produces a factor of $2^{2}=4$. Similarly, $\prod_{p\in\mathfrak{S}}(s^{t}_{p},r^{t}_{p})$ represents the product of the random signs of $s^{t}$ and $r^{t}$ in the Hamiltonian terms. $Z_{c}(E^{c})=\sum_{\{\sigma\}}e^{-H_{3D}(\{\sigma\}|E^{c})}$ is the partition function of a single R4bIM, and $\overline{\cdots}$ denote the thermally averaged value with respect to $H_{3D}(\{\sigma\}|E^{c})$. $\mathfrak{S}\cap(t=T)$ denotes the loop at the intersection between the surface $\mathfrak{S}$ and the time slice $t=T$, see Fig. 9(b). Since the $\{\sigma\}$ spins are in the deconfined phase, we expect that the thermal average for a typical $E^{c}$ follows the perimeter-law, so that $\overline{\prod_{l\in\gamma}\sigma_{l}}\sim e^{-A|\gamma|}$ for any loop $\gamma$ with some constant $A>0$. Therefore, when the $\{\sigma\}$ spins are summed over, the $\{\tau\}$ spins are in an effective R4bIM, whose coupling constant $J$ is modified as
$\tanh J\to\tanh J_{\text{eff}}=e^{-A}\tanh J$ (24)
for all the terms in Eq. (22) that are located near the plane defect. Intuitively, this effect can be regarded as an increase of local temperature near $t=T$. Hence, with increasing error rates $p$ and $q$, the $\{\tau\}$ spins near $t=T$ will enter a local confined phase before the confinement transition for the rest of the system. From the viewpoint of decoding, this implies that the target code block has a finite but lower error threshold near $t=T$ compared to the rest of the system.
For a rough, conservative estimate for the local error threshold of the target block near $t=T$, we consider the situation where the Wilson loop of the $\{\sigma\}$ spin renormalizes the coupling constant $J$ to $J_{\text{eff}}$ in a considerable region around $t=T$. If we limit ourselves to the case where the syndrome error rate $q$ equal the bit-flip rate $p$, setting $J=K$, the target block reaches threshold when $J_{\text{eff}}(p)$ reaches its critical value:
$\tanh J_{\text{eff}}(p^{t}_{\text{c}})=e^{-A(p^{t}_{\text{c}})}(1-2p^{t}_{\text{c}})=\tanh J(p^{*})=1-2p^{*}.$ (25)
This implies that the reduction of error threshold from that of memory with imperfect syndromes, $p^{*}=0.033$ will be moderate as long as the loop tension $A$ is small. The numerical study of R4bIM in Ref. [10] found that $A$ is of order $0.01$ when $p$ is close but smaller than $p^{*}$. Therefore, we estimate that the solution to Eq. (25) by setting $A=0.01$, which yields $p_{\text{c}}^{t}=0.028$. Hence, even with conservative assumptions, the decrease in the error threshold of the target block due to the spread of syndrome error is not significant. We leave the direct numerical simulation of the local error threshold for future work.
V Conclusion and outlook
Figure 10: (a) A noisy transversal Clifford logical circuit between two code blocks $A$ and $B$ consisting of two tCNOT gates and a (fold-)transversal $S$ gate. Ever pair of transversal gates are separated by $\Delta t$ rounds of syndrome error extractions and generic physical error channels $\mathcal{N}$. (b) Sketch of the impact of transversal gates on the stat-mech model that is mapped from the circuit in (a). The $X/Z$ spacetime detectors of the two code blocks $A$ and $B$, denoted by $\{d^{A/B}_{X/Z}\}$, give rise to 4 different set of classical spins in total, which are coupled at the plane defects induced by the logical operations. The propagation of Pauli errors by the transversal gates are marked by red arrows. The impact region $|t-T|<\lambda$ due to increased effective local temperature are marked by darker colors for each plot.
In summary, we studied the error threshold of the tCNOT gate in toric codes under bit-flip noise using the stat-mech mapping approach. We considered two error models, the persistent bit-flip error and bit-flip plus syndrome errors. These error models quantify the reduction of threshold compared to quantum memory due to two distinct error spreading mechanisms by the tCONT gate: the spread of physical bit-flip error and syndrome error. In case of persistent bit-flip error with perfect syndromes, we mapped the decoding problem to a 2D random AT model. The error thresholds of control and target code blocks in this case were $p_{\text{c}}^{c}=0.099$ and $p_{\text{c}}^{t}=0.080$, respectively. These values were rigorously established via Monte-Carlo simulation and FSS. The case with syndrome error was mapped to a plane defect in 3D R4bIM, and the decrease in the error threshold of the target block, $p_{\text{c}}^{t}=0.028$, was conservatively estimated leveraging previously reported numerical results [10]. Comparing these thresholds to the memory thresholds in respective cases, our results demonstrate that both mechanisms of error spreading only cause a minor, non-devastating decrease in the decoding threshold compared to the memory threshold.
Throughout this work, we specifically consider logical circuits with a single tCNOT gate. However, our study can be straightforwardly generalized to map a general transversal Clifford logical circuit of toric codes to a stat-mech model, where each consecutive logical operation is separated by $\Delta t$ rounds of syndrome extractions, see Fig. 10(a). In this stat-mech model of the logical circuit, each transversal Clifford gate operation affects a specific time slice in the 3D stat-mech model. Hence, the effect of each gate at $t=T$ will be local in the time direction inside a time window $|t-T|\leq\lambda$, where $\lambda\sim O(1)$, as shown pictorially in Fig. 10(b). For a given circuit, one should be able to simulate the corresponding models based on the mappings that we have provided in this paper. Therefore, our estimation of error threshold can be readily generalized to estimate the threshold of the entire logical circuit.
Interestingly, our success in mapping transversal logical circuits to stat-mech models signals a potential connection between the error threshold of the logical circuit and topological phase transitions. Namely, the entire noisy logical circuit within and beyond the error threshold can be regarded as two distinct phases of matter. The determination of error threshold of logical circuit can then be formulated as a phase-recognition task suitable for learning agents [52].
Acknowledgments YX thanks Kaavya Sahay and Mark Turner for helpful discussions on tCNOT decoders. YX and E-AK acknowledge support by the NSF through the grant OAC-2118310. YZ and E-AK acknowledge support by the National Science Foundation (Platform for the Accelerated Realization, Analysis, and Discovery of Interface Materials (PARADIM)) under Cooperative Agreement No. DMR-2039380. This research is funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF10436 to E-AK. JPS was funded by NSF DMR 2327094. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
Appendix A Coupling constants in the Hamiltonian of the random Ashkin-Teller model
In this appendix, we derive the coupling constants $K_{2}$ and $K_{4}$ in the random AT model in Eq. (17). To this end, we first rewrite the probability distribution of the random link distribution in Eq. (III.3) in an exponential ansatz:
$p(s^{c},s^{t})\propto e^{As^{c}+Bs^{t}+Cs^{c}s^{t}},$ (26)
where $A,B$ and $C$ are constants to be fixed from the equality. Expanding the exponential using the equality $e^{As}=\cosh A+s\sinh A$ for any binary number $s=\pm 1$, we have
$(1-2\tilde{p})^{2}=\frac{\sinh A\cosh B\cosh C+\cosh A\sinh B\sinh C}{\cosh A\cosh B\cosh C+\sinh A\sinh B\sinh C},$
$(1-2\tilde{p})^{3}=\frac{\cosh A\sinh B\cosh C+\sinh A\cosh B\sinh C}{\cosh A\cosh B\cosh C+\sinh A\sinh B\sinh C}=\frac{\cosh A\cosh B\sinh C+\sinh A\sinh B\cosh C}{\cosh A\cosh B\cosh C+\sinh A\sinh B\sinh C}.$ (27)
Solving the equations above, we have
$A=\frac{1}{4}\log\frac{[1+(1-2\tilde{p})^{2}]^{2}-4(1-2\tilde{p})^{6}}{[1-(1-2\tilde{p})^{2}]^{2}}=K_{2},$
$B=C=\frac{1}{4}\log\frac{1+(1-2\tilde{p})^{2}+2(1-2\tilde{p})^{3}}{1+(1-2\tilde{p})^{2}-2(1-2\tilde{p})^{3}}=K_{4}.$ (28)
In the final equalities of both lines, we redefined the constants to be $K_{2}$ and $K_{4}$, which match the coupling constants in the Hamiltonian in Eq. (17) once we substitute $s^{c/t}_{l}$ according to Eq. (15). Both $K_{2}$ and $K_{4}$ decrease monotonically with increasing $\tilde{p}$ (see Fig. 4), leading to order-disorder phase transitions of the $\{\sigma\}$ and $\{\tau\}$ spins. Since $K_{2}>K_{4}$, the $\{\tau\}$ spins disorders before $\{\sigma\}$ spins with increasing $\tilde{p}$.
Appendix B Derivation of the stat-mech model with syndrome errors
To parameterize the error $E$ with syndrome errors, we use a binary number $r^{c,t}_{p}(t)=1/-1$ to denote that the syndrome extraction of the $Z$ stabilizer at the plaquette $p$ in the control or target block is correct/incorrect, respectively. $t\in\mathbb{Z}$ in the bracket is the time step at which the syndrome extraction happens. The probability distribution is given by $p_{r}(r^{c,t}_{p}(t))=\frac{1+(1-2q)r^{c,t}_{p}(t)}{2}$. The bit-flip errors are now represented by $s_{l}^{c,t}(t+\frac{1}{2})=\pm 1$ with probability $p_{s}(s_{l}^{c,t}(t+\frac{1}{2}))=\frac{1+(1-2p)s_{l}^{c,t}(t+\frac{1}{2})}{2}$. Without the tCNOT gate, the values of the spacetime detectors are given by
$d^{c}(t+\frac{1}{2})=r^{c}_{p}(T)r^{c}_{p}(t+1)\prod_{l\in p}s_{l}^{c}(t+\frac{1}{2})\text{ and }d^{t}(t+\frac{1}{2})=r^{t}_{p}(t)r^{t}_{p}(t+1)\prod_{l\in p}s_{l}^{t}(t+\frac{1}{2}).$ (29)
For a tCNOT gate that happens right after time step $T$, the values of the spacetime detectors of the control and target blocks become
$d^{c}(T+\frac{1}{2})=r^{c}_{p}(T)r^{c}_{p}(T+1)\prod_{l\in p}s_{l}^{c}(T+\frac{1}{2})\text{ and }d^{t}(T+\frac{1}{2})=r^{t}_{p}(T)r^{c}_{p}(T)r^{t}_{p}(T+1)\prod_{l\in p}s_{l}^{t}(T+\frac{1}{2}),$ (30)
respectively. Note that since $r^{c}_{p}(T)$ is in three spacetime detectors, $d^{c}(T-\frac{1}{2})$, $d^{c}(T+\frac{1}{2})$ and $d^{t}(T+\frac{1}{2})$, the decoding graph across the tCNOT gate contains weight-3 hyperedges. Away from $t=T$, $r^{c}_{p}$ do not enter the spacetime detector $d^{t}(t\neq T)$ of the target block.
To obtain the stat-mech model in this case, we need to parameterize the trivial cycles $C$ that lead to the same set of syndromes $\{d^{c},d^{t}\}$. Based on their forms in Eqs. (29) and (30), we introduce two sets of Ising spins $\{\sigma\}$ and $\{\tau\}$ for every link of the cubic lattice. When $t\leq T$, the spacetime detectors are invariant under the following reparameterization via the Ising spins:
$s^{c}_{l}(t-\frac{1}{2})\to s^{c}_{l}(t-\frac{1}{2})\sigma_{l}(t-1)\sigma_{l}(t)\sigma_{l_{1}}(t-\frac{1}{2})\sigma_{l_{2}}(t-\frac{1}{2}),\ s^{t}_{l}(t-\frac{1}{2})\to s^{t}_{l}(t-\frac{1}{2})\tau_{l}(t-1)\tau_{l}(t)\tau_{l_{1}}(t-\frac{1}{2})\tau_{l_{2}}(t-\frac{1}{2}),$
$r^{c}_{p}(t)\to r^{c}_{p}\prod_{l\in p}\sigma_{l}(t),\ r^{t}_{p}(t)\to r^{t}_{p}\prod_{l\in p}\tau_{l}(t),$ (31)
the same as the case without the tCNOT gate. At $t>T$, to account for $r^{c}_{p}(T)$ in the detector $d^{t}(T)$ in Eq. (30), the reparameterization is changed accordingly:
$s^{c}_{l}(t-\frac{1}{2})\to s^{c}_{l}(t-\frac{1}{2})\sigma_{l}(t-1)\sigma_{l}(t)\sigma_{l_{1}}(t-\frac{1}{2})\sigma_{l_{2}}(t-\frac{1}{2}),$
$s^{t}_{l}(t-\frac{1}{2})\to s^{t}_{l}(t-\frac{1}{2})\sigma_{l}(t-1)\tau_{l}(t-1)\sigma_{l}(t)\tau_{l}(t)\sigma_{l_{1}}(t-\frac{1}{2})\tau_{l_{1}}(t-\frac{1}{2})\sigma_{l_{2}}(t-\frac{1}{2})\tau_{l_{2}}(t-\frac{1}{2}),$
$r^{c}_{p}(t)\to r^{c}_{p}(t)\prod_{l\in p}\sigma_{l}(t),\ r^{t}_{p}(t)\to r^{t}_{p}(t)\prod_{l\in p}\sigma_{l}(t)\tau_{l}(t),$ (32)
that is, the Ising spin $\{\tau\}$ becomes $\{\tau\sigma\}$ once we go from $t\leq T$ to $t>T$. In other words, the tCNOT gate induces a domain wall that permutes the Ising spins.
Putting everything together, the probability of having the same syndrome set $\{d^{c}(t),d^{t}(t)\}$ without logical error sums up to
$\sum_{C}\mathrm{prob}((E^{c},E^{t})\cdot C)$
$=\sum_{\{\sigma\},\{\tau\}}\prod_{t\neq T}\left[\prod_{l}p_{s}\left(s^{c}_{l}(t+\frac{1}{2})\sigma_{l}(t)\sigma_{l}(t+1)\sigma_{l_{1}}(t+\frac{1}{2})\sigma_{l_{2}}(t+\frac{1}{2})\right)p_{s}\left(s^{t}_{l}(t+\frac{1}{2})\tau_{l}(t)\tau_{l}(t+1)\tau_{l_{1}}(t+\frac{1}{2})\tau_{l_{2}}(t+\frac{1}{2})\right)\right]$
$\times\prod_{l}p_{s}\left(s^{c}_{l}(T+\frac{1}{2})\sigma_{l}(T)\sigma_{l}(T+1)\sigma_{l_{1}}(T+\frac{1}{2})\sigma_{l_{2}}(T+\frac{1}{2})\right)p_{s}\left(s^{t}_{l}(T+\frac{1}{2})\sigma_{l}(T)\tau_{l}(T)\tau_{l}(T+1)\tau_{l_{1}}(T+\frac{1}{2})\tau_{l_{2}}(T+\frac{1}{2})\right)$
$\times\prod_{t}\left[\prod_{p}p_{r}\left(r^{c}_{p}(t)\prod_{l\in p}\sigma_{l}(t)\right)p_{r}\left(r^{t}_{p}(t)\prod_{l\in p}\tau_{l}(t)\right)\right].$ (33)
Note that we have redefined the spins $\{\tau\sigma\}$ to be $\{\tau\}$ at $t>T$ in the summation. Using $p_{s}(s)\propto e^{Js}$ and $p_{r}(r)\propto e^{Kr}$, where $J=\frac{1}{2}\ln\frac{1-p}{p}$ and $K=\frac{1}{2}\ln\frac{1-q}{q}$, the probability is proportional to the partition function
$\sum_{C\in S}\mathrm{prob}((E^{c},E^{t})\cdot C)\propto Z(E^{c},E^{t})=\sum_{\{\sigma\},\{\tau\}}e^{-H_{3D}(\{\sigma\},\{\tau\}|E^{c},E^{t})}$ (34)
$H_{3D}(\{\sigma\},\{\tau\}|E^{c},E^{t})=-J\sum_{l,t}s^{c}_{l}(t+\frac{1}{2})\sigma_{l}(t)\sigma_{l}(t+1)\sigma_{l_{1}}(t+\frac{1}{2})\sigma_{l_{2}}(t+\frac{1}{2})$
$\quad\quad\quad-J\sum_{l,t\neq T}s^{t}_{l}(t+\frac{1}{2})\tau_{l}(t)\tau_{l}(t+1)\tau_{l_{1}}(t+\frac{1}{2})\tau_{l_{2}}(t+\frac{1}{2})$
$\quad\quad\quad-J\sum_{l}s^{t}_{l}(T+\frac{1}{2})\sigma_{l}(T)\tau_{l}(T)\tau_{l}(T+1)\tau_{l_{1}}(T+\frac{1}{2})\tau_{l_{2}}(T+\frac{1}{2})-K\sum_{l,t}\left[r^{c}_{p}(t)\prod_{l\in p}\sigma_{l}(t)+r^{t}_{p}(t)\prod_{l\in p}\tau_{l}(t)\right].$ (35)
Compared to the Hamiltonian of two decoupled R4bIMs, $H_{3D}(\{\sigma\}|E^{c})+H_{3D}(\{\tau\}|E^{t})$, the difference in the Hamiltonian above is the terms that involve $s^{t}_{l}(T+\frac{1}{2})$, which contain the $\sigma$ spins.